OpenStudy (anonymous):

Using the ratio of perfect squares method, what is square root of 96 rounded to the nearest hundredth? Again with the square root of 14.

OpenStudy (anonymous):

Hi @Ogziii

OpenStudy (anonymous):

1st one is : 4.578 2nd one is : 2.410

OpenStudy (anonymous):

To the nearest hundredths?

OpenStudy (anonymous):

@texaschic101 @geerky42 Can any of you help me? PLZ

OpenStudy (anonymous):

@ogziii I need it to the nearest hundredths and how did you get that?

geerky42 (geerky42):

hmm, well square root of 96 is in between \(\sqrt{81}=9\) and \(\sqrt{100}=10\) and 96 is closer ti 100 than 81

OpenStudy (anonymous):

So we have a ratio of 4/19

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Hello @texaschic101 ... can you help me with this problem?

OpenStudy (texaschic101):

sq rt 96 = sq rt 16 * sq rt 6 = 4 sq rt 6 = 9.798.....rounds to 9.80

OpenStudy (anonymous):

is that for 96 or 14

OpenStudy (texaschic101):

that was for 96.... not sure about sq rt 14 because there is no perfect squares in there the answer will be : 3.74.....but thats from just looking at a calc

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

thx

OpenStudy (texaschic101):

sure thing :)